OpenStudy (anonymous):

solve the absolute value equation graph the solution |3x+1|=1

sam (.sam.):

Square both sides

OpenStudy (anonymous):

so like? _1,1 on a line graph?

sam (.sam.):

?

sam (.sam.):

|3x+1|=1 \[(3x+1)^2=1^2\]

sam (.sam.):

Then solve it find x